# RMT-BVQA: Recurrent Memory Transformer-based Blind Video Quality Assessment for Enhanced Video Content

### v2

@@ 4 Experimental Configuration
Implementation Details. The implementation of the RMViT module is based on the Vision transformer for small dataset [28], where the depth is set to 8 and the number of heads is 64. The hidden dimension is adjusted to 2048 following the practice in [37]. The size of the segment is fixed at 4. The number of memory tokens and the segment length are 12 (this is verified in Supplementary through an analysis of the training results). The batch size for training is $B$ = 256. $\tau$ is set to 0.1. The threshold $\mathrm{TH}$, used in the quality classification, is set to 6 based on [13]. Data augmentation is performed during the training content generation through block rotation. The RMViT module, the prediction network, $f(\cdot)$, and the projector network, $g(\cdot)$, were trained simultaneously from scratch for 150 epochs using stochastic gradient descent (SGD) at a learning rate of 0.00025. A linear warmup over the first ten epochs was applied to the learning rate, followed by a cosine decay schedule used in [33]. The proposed method was implemented using Pytorch 1.13 with an NVIDIA 3090 GPU.
@@
Evaluation settings. To demonstrate the effectiveness of the proposed RMT-BVQA, we compared it with two classic, two regression-based and six deep blind VQA methods. Due to the limited test content available, we performed a five-fold cross validation experiment based on the VDPVE database [^2] Here we refer to the VDPVE training set, which contains 839 sequences, while the test set is not accessible. for RMT-BVQA and all the other eight learning-based VQA methods. This dataset is further divided into three sub-datasets: Subset A including 414 videos generated through colour, brightness, and contrast enhancement; Subset B comprising 210 deshaked videos; and Subset C containing 215 deblurred videos. It should be noted that for the proposed method, only the linear regression is optimised using the training set in the cross validation. The five-fold cross validation has been performed 100 times to calculate the average correlation coefficients. To test the correlation performance with subjective opinions, we employed both the SRCC and the Pearson Linear Correlation Coefficient (PLCC) as evaluation metrics. We have also compared our approach with the other ten VQA methods in terms of their computational complexity. However, due to limited space in the main paper, we include the complexity results in the Supplementary.
5 Experiments and Results
Table 2 summarises the comparison results between our proposed method, RMT-BVQA, and ten benchmark VQA methods [^3] It is noted that the results presented here are different from [32], where the test set (unavailable publicly) of the VDPVE database was employed for evaluation.. It can be observed that RMT-BVQA offers the best overall performance among all the tested quality metrics - the only one with SRCC and PLCC values above 0.8. For all three subsets, A, B, and C in the VDPVE database, which correspond to different enhancement method types, our model is the best performer on Subset A and B, and the second best on Subset C (very close to FAST-VQA). Figure 2 provides three visual examples in which the proposed RMT-BVQA offers the correct quality differentiation as human perception, while the second best performer, FAST-VQA [52] fails to do that.
To further verify the effectiveness of the main contributions in this work, an ablation study is also conducted including the following sub tests.
Recurrent Memory Vision Transformer. We evaluated the contribution of our RMViT module by replacing it with an alternative network structure, Gates Recurrent Unit (GRU) [7], which has been utilised in previous contrastive learning based video quality assessment tasks [38]. The new variant is denoted by (v1-GRU).
@@ 5 Experiments and Results
Figure 2: Visual examples from three subsets of VDPVE [14] (from the left to the right: colour transform, deshaking and deblurring) demonstrating the superiority of the proposed method. In these three cases, RMT-BVQA (ours) provides the same quality differentiation as human perception does.
@@
Content-quality-aware contrastive learning. We further verify the effectiveness of the proposed content-quality-aware loss function by removing the content-aware loss, producing (v2-quality). We did not test the contribution of the quality-aware loss, because only employing content-aware loss leads to unstable training performance.
@@
The training database. Since it is difficult to find an alternative to the proposed training database, which can support the optimisation of the RMViT module, we instead compared v1-GRU to the original CONVIQT model to verify the contribution of the training database. v1-GRU effectively has the same network architecture as CONVIQT - differing only in the use of the new training database to optimise the GRU module.
The results of the ablation study are also provided in Table 1, where both v1-GRU and v2-quality achieve lower average correlation coefficients compared to RMT-BVQA. It can be observed that v1-GRU outperforms CONVIQT, with better overall correlation performance as shown in Table 2, confirming the effectiveness of the developed training database.
@@ 6 Conclusion
In this paper, we propose a blind deep VQA method specifically for enhanced video content based on a novel self-supervised learning training methodology. In this work, we designed a new Recurrent Memory Vision Transformer (RMViT) module to obtain video quality representations, which is optimised through contrastive learning based on a content-quality-aware loss function. A large and diverse training dataset has also been developed containing various types of enhanced video content. The proposed method, RMT-BVQA, has been tested on a video enhancement quality database through five-fold cross validation, and exhibits higher correlation with opinion scores when compared to ten existing no-reference VQA methods. Future work should focus on benchmarking on more diverse enhanced video content.

### v5

@@ 4 Experimental Configuration
Implementation Details. The implementation of the RMViT module is based on the Vision transformer for small datasets [32], where the depth is set to 8 and the number of heads is 64. The hidden dimension is adjusted to 2048 following the practice in [41]. The size of the segment is fixed at 4. The number of memory tokens and the segment length are 12 (this is verified in Ablation Study through an analysis of the training results). The batch size for training is $B$ = 256. $\tau$ is set to 0.1. The threshold $\mathrm{TH}$, used in the quality classification, is set to 6 based on [14]. Data augmentation is performed during the training content generation through block rotation. The RMViT module, the prediction network, $f(\cdot)$, and the projector network, $g(\cdot)$, were trained simultaneously from scratch for 150 epochs using stochastic gradient descent (SGD) at a learning rate of 0.00025. A linear warm-up over the first ten epochs was applied to the learning rate, followed by a cosine decay schedule used in [37]. The proposed method was implemented using Pytorch 1.13 with an NVIDIA 3090 GPU.
Benchmark blind VQA methods. The proposed quality metric is compared with ten blind VQA methods, including two conventional models, NIQE [45], BRISQUE [43], three regression-based methods, V-BLIINDS [51], TLVQA [28] and ChipQA [12], and five deep learning based approaches, BVQA [33], VSFA [34], SimpleVQA [30], CONVIQT [42], FAST-VQA [59] and RanKDVQA-NR [14]. It is noted that some well-performing no-reference VQA models, such as TB-VQA [60], which ranks first in the NTIRE 2023 quality assessment of video enhancement challenge [36], and SB-VQA [23], have not been included in this experiment, as their source codes are not public accessible.
@@
Evaluation settings. Due to the limited test content available, we performed a five-fold cross validation experiment based on the VDPVE database for RMT-BVQA and all the other eight learning-based (regression or deep learning) VQA methods. It should be noted that here we refer to the VDPVE training set, which contains 839 sequences, while the corresponding test set is not accessible, as mentioned in Sec. 2. This dataset is further divided into three sub-datasets: Subset A including 414 videos generated through colour, brightness, and contrast enhancement; Subset B comprising 210 deshaked videos; and Subset C containing 215 deblurred videos. It should be noted that for the proposed method, only the linear regression is optimized using the training set in the cross validation. The five-fold cross validation has been performed 100 times (the split is based on source content) to calculate the average correlation coefficients. All trainable methods (with their pre-trained models) were fine-tuned for up to 150 epochs during each cross validation with an early stopping strategy. To test the correlation performance with subjective opinions, we employed both the SRCC and the Pearson Linear Correlation Coefficient (PLCC) as evaluation metrics. We have also compared the computational complexity of our approach with the other five deep learning-based VQA methods in terms of runtime and model parameters (see Tab. 4).
5 Results and Discussion
@@
5.1 Overall performance
Tab. 2 summarizes the comparison results between our proposed method, RMT-BVQA, and ten benchmark VQA methods. It should be noted that the results presented here are different from [36], where the test set (unavailable publicly, as mentioned in in Sec. 2) of the VDPVE database was employed for evaluation. In this experiment, we performed cross validations (100 times) on the VDPVE training set instead. It can be observed in Tab. 2 that RMT-BVQA offers the best overall performance among all the tested quality metrics - the only one with SRCC and PLCC values above 0.8. For all three subsets, A, B, and C in the VDPVE database, which correspond to different enhancement method types, our model is the best performer on Subset A and B, and the second best on Subset C (very close to FAST-VQA). Fig. 3 provides three visual examples in which the proposed RMT-BVQA offers the correct quality differentiation as human perception, while the second (FAST-VQA [59]) or the third best performer (CONVIQT [42]) fails to do that.
@@ 5.1 Overall performance
Figure 3: Visual examples from three subsets of VDPVE [15] (from the top to the bottom: color transform, deshaking and deblurring) demonstrating the superiority of the proposed method. In these three cases, RMT-BVQA (ours) provides the same quality differentiation as human perception does.
5.2 Ablation study
To further verify the effectiveness of the main contributions in this work, an ablation study is also conducted including the following sub-tests.
Recurrent Memory Vision Transformer. We evaluated the contribution of our RMViT module by replacing it with an alternative network structure, Gated Recurrent Unit (GRU) [8], which has been utilized in previous contrastive learning based video quality assessment tasks [42]. The new variant is denoted by (v1-GRU), and its results (based on the same cross-validation experiment) are shown in Tab. 2. It can be observed that v1-GRU achieves lower average correlation coefficients compared to the full RMT-BVQA, in particular on Subset B (deshaking). This verifies the contribution of the recurrent memory vision transformer, due to its long-term scene memory compared to the traditional GRU module, especially for those videos containing severe scene transitions (in Subset B).
Moreover, to determine the values of two hyperparameters in this RMViT module, including the number of memory tokens, $M$, and the segment length, $S$ (used in training), we performed an analysis based on the training performance by varying these values within predefined ranges (limited by our memory constraints), $2\leq M\leq 12$ and $2\leq S\leq 12$. As shown in Tab. 3, when both the length of segments and the number of memory tokens are set to 12 (we cannot test higher values due to memory constraints), the training loss reaches its lowest level. This justifies the selection of these two hyperparameter values.
Table 3: The analysis on the memory token size, $M$ and the segment length $S$, used in the training process. Here the size of the red circle indicates the value of the lowest training loss.
The number of memory tokens $M$
Loss 2 4 6 8 10 12
$S$ 12 6.236 5.927 5.728 5.475 5.151 5.133
10 6.374 5.762 5.783 5.564 5.174 5.166
8 6.490 6.271 5.949 5.770 5.569 5.492
6 6.698 6.487 6.174 6.052 5.961 5.944
4 6.765 6.519 6.434 6.326 6.309 6.341
2 6.889 6.879 6.747 6.654 6.541 6.553
@@ 5.2 Ablation study
Content-quality-aware contrastive learning. We further verify the effectiveness of the proposed content-quality-aware loss function by removing the content-aware loss, producing (v2-quality). We did not test the contribution of the quality-aware loss, because only employing content-aware loss leads to unstable training performance. The results in Tab. 2 show that v2-quality is worse than the original RMT-BVQA, which confirms the effectiveness of the content-quality-aware contrastive learning methodology.
@@
The training database. Since it is difficult to find an alternative to the proposed training database, which can support the optimization of the RMViT module, we instead compared v1-GRU to the original CONVIQT model to verify the contribution of the training database. v1-GRU effectively has the same network architecture as CONVIQT - differing only in the use of the new training database to optimize the GRU module. From Tab. 2, we can observe that v1-GRU outperforms CONVIQT, with better overall correlation performance - this justifies the contribution of the developed training database.
5.3 Model Complexity
Tab. 4 provides the computational complexity results for the proposed method and the other five deep learning based benchmark methods, in terms of runtime (second, for processing 300 frames 1280$\times$720 video) and the number of parameters. This experiment is implemented on a workstation with an NVIDIA 3090 GPU, a 3.30 GHz Intel W-1250 CPU and 64 GB of RAM. The high complexity in particular in model size is mainly due to the complex network structure employed (vision transformer in the RMViT module). Although our parameters are significantly higher compared to the benchmarks, the runtime has not increased proportionally, showing only a 30.48% increase compared to CONVIQT [42]. This discrepancy arises from considering only the memory tokens generated by the final two segments during inference, thereby substantially reducing the model’s effective runtime.
Table 4: The runtime and model size (number of parameters) figures for the proposed method and the other learning-based benchmarks.
Benchmark     Runtime (s)       Params (M)
BVQA [33] 13.9 57.1
VSFA [34] 6.4 0.5
CONVIQT [42] 18.7 38.6
       FAST-VQA [59] 1.4 27.6
RankDVQA-NR [14] 46.2 4.6
RMT-BVQA 24.4 703.7
6 Limitations of the proposed method
Although our proposed method shows superior performance over the other benchmarked quality metrics, it is also associated with higher complexity, in particular in model size (Tab. 4). This could potentially lead to greater energy consumption and a larger carbon footprint when implemented in practical applications. This issue can be alleviated by integrating dynamic input token pruning [50, 56, 27], into the RMViT module. Approaches such as knowledge distillation [21] and model compression [2] can also be applied to further reduce the complexity of the model while maintaining performance. Moreover, the temporal pooling method employed in RMT-BVQA can be improved through more advanced feature fusion models [19].
@@ 7 Conclusion
In this paper, we propose a deep blind VQA method specifically for enhanced video content based on a novel self-supervised learning training methodology. In this work, we designed a new Recurrent Memory Vision Transformer (RMViT) module to obtain video quality representations, which is optimized through contrastive learning based on a content-quality-aware loss function. A large and diverse training dataset has also been developed containing various types of enhanced video content, which supports the proposed contrastive learning methodology but does not rely on expensive subjective tests to obtain ground-truth quality labels. The proposed method, RMT-BVQA, has been tested on a video enhancement quality database through five-fold cross validation, and exhibits higher correlation with opinion scores when compared to ten existing no-reference VQA methods. Future work should focus on benchmarking (and training) on more diverse enhanced video content and reducing the high complexity in particular in model size (Tab. 4). The prediction performance of the proposed method on deblurred content (Subset C) can also be further improved.
Acknowledgements
Research reported in this paper was supported by an Amazon Research Award, Fall 2022 CFP. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of Amazon. The authors also appreciate the funding from the UKRI MyWorld Strength in Places Programme (SIPF00006/1).